[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed May 17, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed May 17 18:45:33 UTC 2017


HSU Shift Log for Wed May 17 2017 14:45:27 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: GOES-16 Full Disk imagery for NAWIPS is only coming in  
at :00 and :15. The :30 and :45 images are missing. CONUS imagery appears  
to be OK.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170517/9f05219e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list